OpenStudy (anonymous):

Solve the system of equations by graphing. 2x-y=2 , 5x+y=5

OpenStudy (anonymous):

2x-y=2 5x+y=5 put these into y=mx+b form

OpenStudy (anonymous):

and you should get

OpenStudy (anonymous):

and for the graph

OpenStudy (anonymous):

and when you put the equation5x+y=5in y=mx+b form you get

OpenStudy (anonymous):

do you know the plotting points for each equation? @Tahmina

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so the solution is y=0 x=1

OpenStudy (anonymous):

y= 2x-2 when x=0 y=2 plot this

OpenStudy (anonymous):

now i feel like im doing your hw

OpenStudy (anonymous):

No i mean the plotting points on the graph